You would like to put a GEO satellite for QU_SAT in orbit. Its weight is 3000 Kg. What should be its

exact altitude for being stationary with respect to Qatar. Show the steps of your calculations.
Answer #1

A satellite at altitude h from earth surface is at distance R+h from center of earth where R is radius of earth.

According to Newton's law of universal gravitation, the force of attraction between a satellite with mass m at altitude h is:

F = G\frac{Mm}{(R+h)^2}

Where G = gravitational constant = 6.674x10-11 m3/(kgs2)
            R = radius of earth = 6 378 km
            M = mass of earth = 5.972 × 1024 kg
            m = mass of satellite = 3000 kg, as given in question

A geostationary satellite orbits round earth in 24 hours, thus effectively being stationary w.r.t a certain place on earth surface.

=> The satellite moves through an angle of 2\pi rad (assuming circular orbit for simplicity of calculation) in 24x3600 seconds

=> angular velocity of satellite:

\omega = \frac{2\pi}{24\times 3600}

The centripetal force required to keep the satellite in orbit is given by the formula

F = m(R+h)\omega^2

When a satellite is in orbit the gravitational force must equal the centripetal force, i.e.

G\frac{Mm}{(R+h)^2} = m(R+h)\omega^2

\Rightarrow R+h = \left(\frac{GM}{\omega^2} \right )^{\frac{1}{3}}
Putting the values of G, M and \omega in above equation,

R+h \approx 422.4x105 m = 42240 km

=> h = (42240 - 6378) km = 35862 km [Ans]

We find that altitude of geostationary satellite is independent of it's mass.